### v2

@@ Abstract
As networks continue to expand and become more interconnected, the need for novel malware detection methods becomes more pronounced. Traditional security measures are increasingly inadequate against the sophistication of modern cyber attacks. Deep Packet Inspection (DPI) has been pivotal in enhancing network security, offering an in-depth analysis of network traffic that surpasses conventional monitoring techniques. DPI not only examines the metadata of network packets, but also dives into the actual content being carried within the packet payloads, providing a comprehensive view of the data flowing through networks. While the integration of advanced deep learning techniques with DPI has introduced modern methodologies into malware detection and network traffic classification, state-of-the-art supervised learning approaches are limited by their reliance on large amounts of annotated data and their inability to generalize to novel, unseen malware threats. To address these limitations, this paper leverages the recent advancements in self-supervised learning (SSL) and few-shot learning (FSL). Our proposed self-supervised approach trains a transformer via SSL to learn the embedding of packet content, including payload, from vast amounts of unlabeled data by masking portions of packets, leading to a learned representation that generalizes to various downstream tasks. Once the representation is extracted from the packets, they are used to train a malware detection algorithm. The representation obtained from the transformer is then used to adapt the malware detector to novel types of attacks using few-shot learning approaches. Our experimental results demonstrate that our method achieves classification accuracies of up to 94.76% on the UNSW-NB15 dataset and 83.25% on the CIC-IoT23 dataset.
{IEEEkeywords}
Malware Detection, Malware Classification, Deep Packet Inspection, Self-Supervised Learning, Few-Shot Learning
1 Introduction
@@
With the continuing expansion of the digital landscape, the significance of malware detection becomes increasingly important. With the modernization of sophisticated cyber attacks, traditional security measures often fall short in providing adequate protection. Traditional firewalls and antivirus software often prove inadequate because they primarily rely on predefined packet header rules through conventional packet filtering and known malware signatures, leaving them vulnerable to new and sophisticated cyber attacks that do not match any existing patterns or behaviors [1, 2]. Deep Packet Inspection (DPI) is a well-known tool for the analysis of network packets in the security of networks [3, 4]. DPI goes beyond conventional network monitoring tools by examining not only the five-tuple of network packets, such as source and destination Internet Protocol (IP) addresses, but also inspects the data payload within each packet. This in-depth analysis allows for a more comprehensive understanding of the nature of the traffic passing through a network, enabling the identification of malicious activities that might otherwise go unnoticed.
With the recent advancements in deep learning and artificial intelligence [5, 6, 7], the integration of these fields into network traffic analysis and DPI have become more prevalent. The dynamic nature of cyber attacks, characterized by constantly evolving malware, necessitates a shift from traditional, signature-based detection methods to more adaptive and intelligent systems. It is estimated that 560,000 new pieces of malware are detected daily and more than 1 billion malware programs exist [8]. Deep learning, with its capability to model complex patterns and adapt to evolving information through neural networks, presents a promising approach to addressing the challenges of modern malware detection. With this, cybersecurity systems can analyze data in real time, identifying threats based on behavioral patterns rather than relying solely on known signatures.
Table 1: Comparative Summary of Related Works and the Proposed Approach
@@
Proposed Method Network and IoT Intrusions Raw Packet Bytes Self-Supervised Transformer + FSL Learns embeddings from raw payload; adapts quickly to novel/unseen attacks with few samples.
DPI faces several critical challenges that limit its efficiency and effectiveness. First, most existing methods focus primarily either on statistical-based features or inspecting a few first bytes of raw packet payload data, restricting their ability to capture deeper insights. Second, supervised learning approaches struggle to generalize beyond known attack types, as they learn representations tailored to the labeled classes in the training data [21, 22, 14], resulting in poor performance when faced with novel threats or emerging attack patterns. Third, these methods heavily rely on large amounts of annotated data, excelling in detecting malware types with abundant labeled samples but performing poorly on malware with limited or no annotations.
This paper introduces a groundbreaking approach to malware detection and classification, overcoming existing limitations by harnessing the power of transformers trained with Self-Supervised Learning (SSL). By leveraging few-shot learning techniques, we enable these models to adapt seamlessly to novel, unseen malware threats, setting a new standard in malware detection. The key contributions of this paper are as follows:
@@
  • Generalization to Unseen Malware Types: Self-Supervised Learning (SSL) is a paradigm that empowers models to extract deep, meaningful representations from vast amounts of unlabeled data–eliminating the need for costly and labor-intensive manual annotation [5, 23, 24]. By combining the transformers’ context-aware learning with the power of SSL, we train models on massive unannotated payload datasets, equipping them to decode intricate patterns within both benign and malicious payloads, enabling them to identify complex patterns across large-scale benign and malicious payloads and achieve generalization to unseen malware types.
  • Learning Unseen Malware Patterns from a few Annotated Payloads: Traditional supervised learning-based payload inspection methods rely on vast amounts of annotated data for each malware type. The challenge becomes even more severe when novel malware variants appear, as limited samples make it nearly impossible to train conventional models effectively. While transformers, pre-trained on unannotated data, offer promising generalization capabilities, they inherently lack the ability to recognize unseen malware types. To bridge this gap, we integrate few-shot learning techniques, enabling the model to rapidly adapt to both previously encountered and entirely novel malware strains using minimal labeled examples. This approach dramatically enhances detection efficiency and accuracy, allowing AI-driven cybersecurity systems to identify and neutralize emerging threats—a crucial advancement in the fight against ever-evolving cyber attacks.
The structure of this paper is as follows: Section 2 reviews related work, while Section 3 provides background on network packets, self-supervised learning, and few-shot learning. Section 4 describes the datasets and pre-processing steps. Section 5 details the model detection assumptions. Section 6 details the proposed model architecture and training protocols. Section 8 presents the evaluation and ablation results, and Section 9 discusses experimental results using AES and Fernet encryption. Section 10 provides further discussion, and Section 11 concludes the paper.
2 Related Work
In this section, we discuss related works employing modern techniques for detecting and mitigating cyber threats, focusing on approaches that leverage machine learning, deep learning, attention mechanisms, and FSL to improve accuracy in various detection tasks, summarized in Table 1.
2.1 Machine/Deep Learning Attack Detection
@@
2.2 Attention-based Attack Detection
@@
2.3 Few-Shot Learning Attack Detection
FSL’s aim is to learn from limited amounts of labeled data. This is particularly important for intrusion detection, where obtaining large, labeled datasets for new or emerging attack types is difficult. The cyber domain has began to implement FSL techniques to reduce the need for labeled datasets. For example, [18] proposed a FSL method to address the data imbalance issue in network intrusion detection by extracting statistical-based network flow features using a CNN and performing FSL analysis downstream. Similarly, researchers implemented a Siamese CNN to extract features for anomaly detection in industrial cyber-physical systems [20]. The CNN features were once again used for FSL downstream to measure similarities between samples, enabling effective detection of anomalies with minimal labeled data. Lastly, another proposed FSL-based intrusion detection model tailored for IoT environments addressed the challenge of limited labeled data and emerging attack types like zero-day attacks [19]. Although these works produce promising results, they primarily rely on simple CNN-based feature extractors. Simple CNN-based extractors often struggle to capture long-term dependencies and subtle variations within packets, limiting their ability to generalize to unseen attack types effectively. SSL addresses these shortcomings by learning robust and generalizable feature representations based off attention. Additionally, prior FSL methods simulate unseen attacks by re-splitting known classes into different training and testing subsets. The model still has exposure to all of the classes distributions, just with limited examples from each. In contrast, our approach completely excludes one of the attack classes during SSL pre-training and then evaluates the model’s ability to detect the unseen class along with the seen classes in the FSL evaluation step. This approach ensures that the untrained class remains completely unseen during the training phase.
3 Background
@@
3.1 Fundamentals of Network Packets
@@
Network packets play a critical role in the transmission of data across networks. These packets contain information for routing and delivery, including the source and destination IP addresses. IP addresses identify the sending and receiving devices, and the source and destination ports, which specify the sending and receiving applications. These elements form what is known as the five-tuple: Source IP Address, Destination IP Address, Source Port, Destination Port, and Transport Protocol. The Transport Protocol component of the five-tuple specifies the type of protocol used for the connection. Transmission Control Protocol (TCP) and User Datagram Protocol (UDP), being the two most common transport layer protocols [21], are the focus in this study. TCP is known for its reliability and ordered delivery of packets, establishing a connection-oriented session between communicating devices. TCP guarantees that data is received exactly as it is sent, making it ideal for applications where accuracy is necessary, such as web browsing and email transmission [28, 29]. UDP is a connectionless protocol, known for its low-latency and efficient data transfer, without the same reliability guarantees as TCP. This is commonly used in streaming media, online gaming, and real-time communications where speed is more crucial than perfect delivery [28, 29]. These two protocols account for the majority of network traffic on the transport layer.
@@
Given that the five-tuple function is to identify connections, it’s important to note that packets also encompass network payloads. These payloads carry the substantive content being transmitted, including text, images, and files, from the source to the destination. Malicious actors often embed malware within these payloads, exploiting the data transfer process to compromise targeted systems [21, 22, 12, 30]. In a given network packet, the maximum payload size is up to 1460 bytes. Payloads that exceed this limit are fragmented to ensure the complete transmission of data across the network, preventing any portion of the content from being lost during transit. Notably, the fragmentation of network payloads depends on the capacity of the network segment, and these fragments can be reassembled either at the endpoint or by intermediate devices. Furthermore, transformer-based architectures are well-suited for processing fragmented payloads since the self-attention mechanism allows the model to capture dependencies across separate fragments, therefore integrating information that is dispersed over multiple packets. Initially, raw packet and payload data is in binary form and then is converted to hexadecimal format for various purposes. The hexadecimal representation offers a human-readable form, simplifying the process of data inspection and analysis for network experts. Fig. 1 shows a basic diagram of the header and payload bytes with a Maximum Transmission Unit (MTU) of 1500 bytes. The MTU represents the largest size of an Ethernet packet or frame that can be transmitted over a network. Furthemore, it is important to note that DPI operates under the principle that the majority of data transmissions occur without encryption, which enables direct inspection of packet content.[31, 32]
@@
3.2 Self-Supervised Learning
@@
SSL learns a representation of the data via a pretext task. The pretext task is defined depending on the application and the type of the dataset. For example, the pretext task in natural language processing (NLP), which involves learning pattern from a sequential dataset (i.e. sentences), is predicting masked words in a sentence. Through prediction of the masked words, the internal transformer architecture learns to extract distributed representations of whole sentences without the masked word, such that the representation carries enough information to predict what the missing word is. Operating on the principle that the data’s structure can offer insights in an unsupervised manner, it seeks to reduce dependence on labeled datasets and learn to recover whole or parts of the features of its original input [23, 24]. MLM is a key component in SSL and introduced a novel methodology to utilize unlabeled data [5]. MLM involves predicting a masked input within a sequence, allowing the model to consider context from both the left and right sides of any masked input. The bidirectional attention grants the model complete access to all surrounding tokens, enhancing its ability to understand the full context of the sequence [5]. We extend this concept to understand the relationship between the bytes in a packet sequence based on the surrounding context. This approach enhances the model’s ability to generalize from the unlabeled data to unseen examples, improving its predictive accuracy on the raw packet bytes
@@
As shown in Fig. 2, the MLM process begins with masked input values, where specific positions (i.e. [MASK]) are hidden from the model. The input values are passed through an embedding layer combined with positional encoding to preserve order sequence. The transformer encoder then processes the masked sequence through self-attention and feed-forward networks, followed by layer normalization. The resulting output embeddings are sent to a masked prediction layer, which predicts the missing values based on the learned contextual representations. The predicted values are compared with the actual masked values using a cost function, and the resulting loss is backpropagated through the model to update all parameters. This modeling approach is effective for tasks demanding a global understanding of an entire sequence’s context.
3.3 Few-Shot Learning
@@
Similar to how humans are able to recognize objects after observing them once, FSL is a specialized approach designed to enable models to learn and generalize from very limited amounts of labeled data [33, 34, 35]. In FSL, a classifier needs to accommodate for new classes not seen in training, given only few examples from each of those classes. One-shot learning is a form of few-shot learning where the model needs to learn given only one example per class [33]. These algorithms must be able to extract rich contextual understanding from the limited amount of seen examples to generalize to new and unseen instances.
@@
Snell et al. [34] introduced Prototypical Networks for FSL. The main idea behind this research was to show that data points belonging to the same class should be close in distance to each other in an embedding space. The authors suggested that each class should have a single representative or prototype, and this prototype is the mean of all data points in that class once they’ve been projected into a shared embedding space. Support and query sets are then used to simulate the few-shot learning scenario during training and testing. The support set consists of a small number of labeled examples for each class and are used to define the prototype by computing the mean vector of the embedded points for each class. The query set contains examples that the model needs to classify. These examples are different from those in the support set and are used to evaluate the performance of the model. Once the query and support sets are embedded into the space, the model computes the Euclidean distance between each query point and the prototypes. The query points are then classified based on the nearest prototype in the space.
4 Data

### v3

@@ Abstract
As networks continue to expand and become more interconnected, the need for novel malware detection methods becomes more pronounced. Traditional security measures are increasingly inadequate against the sophistication of modern cyber attacks. Deep Packet Inspection (DPI) has been pivotal in enhancing network security, offering an in-depth analysis of network traffic that surpasses conventional monitoring techniques. DPI not only examines the metadata of network packets, but also dives into the actual content being carried within the packet payloads, providing a comprehensive view of the data flowing through networks. While the integration of advanced deep learning techniques with DPI has introduced modern methodologies into malware detection and network traffic classification, state-of-the-art supervised learning approaches are limited by their reliance on large amounts of annotated data and their inability to generalize to novel, unseen malware threats. To address these limitations, this paper leverages the recent advancements in self-supervised learning (SSL) and few-shot learning (FSL). Our proposed self-supervised approach trains a transformer via SSL to learn the embeddings of packet content, including payload, from vast amounts of unlabeled data by masking portions of packets, leading to a learned representation that generalizes to various downstream tasks. Once the representation is extracted from the packets, they are used to train a malware detection algorithm. The representation obtained from the transformer is then used to adapt the malware detector to novel types of attacks using few-shot learning approaches. Our experimental results demonstrate that our method achieves classification accuracies of up to 94.76% on the UNSW-NB15 dataset and 83.25% on the CIC-IoT23 dataset.
Index Terms:
Deep Packet Inspection, Few-Shot Learning, Malware Classification, Malware Detection, Self-Supervised Learning
@@
=-21pt
I Introduction
With the continuing expansion of the digital landscape, the significance of malware detection becomes increasingly important. With the modernization of sophisticated cyber attacks, traditional security measures often fall short in providing adequate protection. Traditional firewalls and antivirus software often prove inadequate since they primarily rely on predefined packet header rules through conventional packet filtering and known malware signatures, leaving them vulnerable to new and sophisticated cyber attacks that do not match existing patterns or behaviors [1, 2]. Deep Packet Inspection (DPI) is a well-known tool for the analysis of network packets in the security of networks [3, 4]. DPI goes beyond conventional network monitoring tools by examining not only the five-tuple of network packets, such as source and destination Internet Protocol (IP) addresses, but also inspects the data payload within each packet. This in-depth analysis allows for a more comprehensive understanding of the nature of the traffic passing through a network, enabling the identification of malicious activities that might otherwise go unnoticed.
With recent advances in deep learning and artificial intelligence [5, 6, 7], the integration of these techniques into network traffic analysis and DPI has become more prevalent. The dynamic nature of cyber attacks, characterized by constantly evolving malware, necessitates a shift from traditional, signature-based detection methods to more adaptive and intelligent systems. It is estimated that 560,000 new pieces of malware are detected daily and more than 1 billion malware programs exist [8]. Deep learning, with its capability to model complex patterns and adapt to evolving information through neural networks, presents a promising approach to addressing the challenges of modern malware detection. With this, cybersecurity systems can analyze data in real time, identifying threats based on behavioral patterns rather than relying solely on known signatures.
TABLE I: Comparative Summary of Related Works and the Proposed Approach
@@ I Introduction
Stein et al. [21] Network and IoT Intrusions Raw Packet Bytes Supervised Transformer + FSL Generalizes well in FSL tasks but assumes there is enough labeled data for supervised pre-training.
Proposed Method Network and IoT Intrusions Raw Packet Bytes Self-Supervised Transformer + FSL Learns embeddings from raw unlabeled packet data; adapts quickly to novel/unseen attacks with few samples.
DPI faces several critical challenges that limit its efficiency and effectiveness. First, most existing methods focus primarily on statistical-based features or inspecting the few first bytes of raw packet payload data, restricting their ability to capture deeper insights. Second, supervised learning approaches struggle to generalize beyond known attack types, as they learn representations tailored to the classes labeled in the training data [22, 23, 14], resulting in poor performance when faced with novel threats or emerging attack patterns. Third, these methods heavily rely on large amounts of annotated data, excelling in detecting malware types with abundant labeled samples but performing poorly on malware with limited or no annotations.
By leveraging few-shot learning techniques, we enable these models to adapt seamlessly to novel, unseen malware threats, setting a new standard in malware detection. The key contributions of this paper are as follows:
@@
  • Generalization to Unseen Malware Types: Self-Supervised Learning (SSL) is a paradigm that empowers models to extract deep, meaningful representations from vast amounts of unlabeled data–eliminating the need for costly and labor-intensive manual annotation [5, 24, 25]. By combining the transformers’ context-aware learning with the power of SSL, we train models on massive unannotated payload datasets, equipping them to decode intricate patterns within both benign and malicious payloads. This enables the model to identify complex patterns across large-scale benign and malicious payloads and achieve generalization to unseen malware types.
  • Learning Unseen Malware Patterns from a few Annotated Payloads: Traditional supervised learning-based payload inspection methods rely on vast amounts of annotated data for each malware type [22, 15, 14]. The challenge becomes even more severe when novel malware variants appear, as limited samples make it nearly impossible to train conventional models effectively. While transformers, pre-trained on unannotated data, offer promising generalization capabilities, they inherently lack the ability to recognize unseen malware types. To bridge this gap, we integrate few-shot learning techniques, enabling the model to rapidly adapt to both previously encountered and entirely novel malware strains using minimal labeled examples. This approach dramatically enhances detection efficiency and accuracy, allowing AI-driven cybersecurity systems to identify and neutralize emerging threats—a crucial advancement in the fight against ever-evolving cyber attacks.
The structure of this paper is as follows: Section II reviews related work, while Section III provides background on network packets, self-supervised learning, and few-shot learning. Section IV describes the datasets and pre-processing steps. Section V details the model detection assumptions. Section VI details the proposed model architecture and training protocols. Section VIII presents the evaluation and ablation results, and Section X discusses experimental results using AES and Fernet encryption. Section XI provides further discussion, and Section XII concludes the paper.
II Related Work
In this section, we discuss related works employing modern techniques for detecting and mitigating cyber threats, focusing on approaches that leverage machine learning, deep learning, attention mechanisms, and FSL to improve accuracy in various detection tasks, summarized in Table I.
II-A Machine/Deep Learning Attack Detection
@@
II-B Attention-based Attack Detection
@@
II-C Few-Shot Learning Attack Detection
FSL’s aim is to learn from limited amounts of labeled data. This is particularly important for intrusion detection, where obtaining large, labeled datasets for new or emerging attack types is difficult. The cyber domain has began to implement FSL techniques to reduce the need for labeled datasets. For example, a FSL method was proposed to address the data imbalance issue in network intrusion detection by extracting statistical-based network flow features using a CNN and performing FSL analysis downstream [18]. Similarly, researchers implemented a Siamese CNN to extract features for anomaly detection in industrial cyber-physical systems [20]. The CNN features were once again used for FSL downstream to measure similarities between samples, enabling effective detection of anomalies with minimal labeled data. Lastly, another proposed FSL-based intrusion detection model tailored for IoT environments addressed the challenge of limited labeled data and emerging, zero-day attacks [19].
Although these works produce promising results, they primarily rely on simple CNN-based feature extractors. Simple CNN-based extractors often struggle to capture long-term dependencies and subtle variations within packets, limiting their ability to generalize to unseen attack types effectively. SSL addresses these shortcomings by learning robust and generalizable feature representations based off attention. Additionally, prior FSL methods simulate unseen attacks by re-splitting known classes into different training and testing subsets. The model still has exposure to all of the classes distributions, just with limited examples from each. In contrast, our approach completely excludes one of the attack classes during SSL pre-training and then evaluates the model’s ability to detect the unseen class along with the seen classes in the FSL evaluation step. This approach is similar to our previously published work [21], where one class is explicitly excluded from training to simulate the unseen scenarios. However, in our previous work we relied on supervised learning to pre-train the LLM on scratch from known malware types. Our work in this manuscript differs since we employ SSL techniques to pre-train the transformer on unlabeled data. This helps demonstrate that SSL can learn meaningful representations from the packet data without relying labeled data, a crucial step in real world applications. Overall, this approach ensures that the untrained class remains completely unseen during the training phase.
III Background
@@
III-A Fundamentals of Network Packets
@@
Network packets play a critical role in the transmission of data across networks. These packets contain information for routing and delivery, including the source and destination IP addresses. IP addresses identify the sending and receiving devices, and the source and destination ports, which specify the sending and receiving applications. These elements form what is known as the five-tuple: Source IP Address, Destination IP Address, Source Port, Destination Port, and Transport Protocol. The Transport Protocol component of the five-tuple specifies the type of protocol used for the connection. Transmission Control Protocol (TCP) and User Datagram Protocol (UDP), being the two most common transport layer protocols [22], are the focus in this study. TCP is known for its reliability and ordered delivery of packets, establishing a connection-oriented session between communicating devices. TCP guarantees that data is received exactly as it is sent, making it ideal for applications where accuracy is necessary, such as web browsing and email transmission [28, 29]. UDP, on the other hand, is a connectionless protocol known for its low-latency and efficient data transfer, without the same reliability guaranties as TCP. This is commonly used in streaming media, online gaming, and real-time communications where speed is more crucial than perfect delivery [28, 29]. These two protocols account for the majority of network traffic at the transport layer.
@@
Given that the five-tuple function is to identify connections, it is important to note that packets also encompass network payloads. These payloads carry the substantive content being transmitted, including text, images, and files, from the source to the destination. Malicious actors often embed malware within these payloads, exploiting the data transfer process to compromise targeted systems [22, 23, 12, 30]. In a given network packet, the maximum payload size is up to 1460 bytes. Payloads that exceed this limit are fragmented to ensure the complete transmission of data across the network, preventing any portion of the content from being lost during transit. Notably, the fragmentation of network payloads depends on the capacity of the network segment, and these fragments can be reassembled either at the endpoint or by intermediate devices. Furthermore, transformer-based architectures are well-suited for processing fragmented payloads since the self-attention mechanism allows the model to capture dependencies across separate fragments, therefore integrating information that is dispersed over multiple packets. Initially, raw packet and payload data exists in binary form and is then converted to hexadecimal format for various purposes. The hexadecimal representation offers a human-readable form, simplifying the process of data inspection and analysis for network experts. Fig. 1 shows a basic diagram of the header and payload bytes with a Maximum Transmission Unit (MTU) of 1500 bytes. The MTU represents the largest size of an Ethernet packet or frame that can be transmitted over a network. Furthermore, it is important to note that DPI operates under the principle that the majority of data transmissions occur without encryption, which enables direct inspection of packet content[31, 32].
@@
III-B Self-Supervised Learning
@@
SSL learns a representation of the data via a pretext task. The pretext task is defined depending on the application and the type of the dataset. For example, the pretext task in natural language processing (NLP), which involves learning patterns from a sequential dataset (i.e. sentences), is predicting masked words in a sentence. Through prediction of the masked words, the internal transformer architecture learns to extract distributed representations of entire sentences without the masked word, such that the representation carries enough information to predict what the missing word is. Operating on the principle that the structure of the data can provide insights in an unsupervised manner, it seeks to reduce dependence on labeled datasets and learn to recover whole or parts of the features of its original input [24, 25]. Masked language modeling (MLM) is a key component in SSL and introduced a novel methodology to utilize unlabeled data [5]. MLM involves predicting a masked input within a sequence, allowing the model to consider context from both the left and right sides of any masked input. The bidirectional attention grants the model complete access to all surrounding tokens, enhancing the ability to understand the full context of the sequence [5]. We extend this concept to understand the relationship between the bytes in a packet sequence based on the surrounding context. This approach enhances the model’s ability to generalize from the unlabeled data to unseen examples, improving its predictive accuracy on the raw packet bytes
@@
As shown in Fig. 2, the MLM process begins with masked input values, where specific positions (i.e. [MASK]) are hidden from the model. The input values are passed through an embedding layer combined with positional encoding to preserve the order sequence. The transformer encoder then processes the masked sequence through self-attention and feed-forward networks, followed by layer normalization. The resulting output embeddings are sent to a masked prediction layer, which predicts the missing values based on the learned contextual representations. The predicted values are compared with the actual masked values using a cost function, and the resulting loss is backpropagated through the model to update all parameters. This modeling approach is effective for tasks demanding a global understanding of an entire sequence’s context.
III-C Few-Shot Learning
@@
Similar to how humans are able to recognize objects after observing them once, FSL is a specialized approach designed to enable models to learn and generalize from very limited amounts of labeled data [33, 34, 35]. In FSL, a classifier needs to accommodate for new classes not seen in training, given only few examples from each of those classes. One-shot learning is a form of few-shot learning where the model needs to learn given only one example per class [33, 36]. These algorithms must be able to extract rich contextual understanding from the limited amount of seen examples to generalize to new and unseen instances.
@@
Snell et al. [34] introduced Prototypical Networks for FSL. The main idea behind this research was to show that data points belonging to the same class should be close in distance to each other in an embedding space. The authors suggested that each class should have a single representative or prototype, and this prototype is the mean of all data points in that class once they have been projected into a shared embedding space. Support and query sets are then used to simulate the few-shot learning scenario during training and testing. The support set consists of a small number of labeled examples for each class and are used to define the prototype by computing the mean vector of the embedded points for each class. The query set contains unlabeled examples that the model needs to classify. These examples are different from those in the support set and are used to evaluate the performance of the model. Once the query and support sets are embedded into the space, the model computes the Euclidean distance between each query point and the prototypes. The query points are then classified based on the nearest prototype in the space.
IV Data